A 0.22 kg piece of silly putty is thrown at a 1.2 kg wooden panel that is initially at the rest. The velocity of the silly putty before the collision was 18.5 m/s. The silly putty sticks to the wooden panel and they move as one after the collision. What is the velocity of the combined objects after the collision.
1
Expert's answer
2020-11-05T07:00:22-0500

We can find the velocity of the combined objects after the collision from the Law of Conservation of Momentum:


"m_pv_p+m_wv_w=(m_p+m_w)v,"

here, "m_p=0.22\ kg" is the mass of a piece of putty, "m_w=1.2\ kg" is the mass of a wooden panel, "v_p=18.5\ \dfrac{m}{s}" is the velocity of the piece of putty before the collision, "v_w=0\ \dfrac{m}{s}" is the velocity of the wooden panel before the collision, "v" is the velocity of the combined objects after the collision.

Then, from this equation we can find "v":


"v=\dfrac{m_pv_p}{m_p+m_w}=\dfrac{0.22\ kg\cdot 18.5\ \dfrac{m}{s}}{0.22\ kg+1.2\ kg}=2.86\ \dfrac{m}{s}."

Answer:

"v=2.86\ \dfrac{m}{s}."